OpenStudy (christopher1448):

HELP WILL GIVE MEDAL!! WILL GIVE MEDAL!! (Answered 1st/ 2nd part Part i did) Part 1. At her roadside stand, Mrs. Farmer displayed her 32 jars of jam by flavor. If only 3/8 of her jams are strawberry, how many strawberry jam does she have? (answer is 12) Part 2. Another 3/8 of her jars of jam are grape. All the rest are blueberry jam. How many jars of blueberry jam does Mrs. Farmer have? Answer( 8 jars) 3/4 of her jars of jam have red labels and the rest have green labels how many jars of jam have red labels??

OpenStudy (christopher1448):

Help!!!

OpenStudy (catpidigm):

32/4=8, 8*3=24, 32-24=8, So it's 8 there you go

OpenStudy (christopher1448):

Thanks!!

OpenStudy (catpidigm):

:)

OpenStudy (catpidigm):

np